### v2

# GraphGPT: Generative Pre-trained Graph Eulerian Transformer

Abstract
We introduce GraphGPT, a novel self-supervised generative pre-trained model for graph learning based on the Graph Eulerian Transformer (GET). First, we propose GET, which combines a standard transformer encoder or decoder architecture with an innovative graph-to-sequence transformation method. This method converts graphs or sampled subgraphs into sequences of tokens representing nodes, edges, and attributes in a reversible manner using Eulerian paths. We pre-train GET using either of the two self-supervised tasks: next-token prediction (NTP) and scheduled masked-token prediction (SMTP). The pre-trained model is then fine-tuned for downstream tasks such as graph-, edge-, and node-level prediction. Despite its simplicity, GraphGPT achieves performance comparable to or surpassing state-of-the-art methods on multiple large-scale Open Graph Benchmark (OGB) datasets. It demonstrates exceptional results on the molecular property prediction dataset PCQM4Mv2 and the protein-protein interaction dataset ogbl-ppa. Notably, generative pre-training enables scaling GraphGPT to 2 billion parameters while maintaining performance gains — a breakthrough that overcomes the scalability limitations of traditional Graph Neural Networks (GNNs) and prior graph transformers (GTs). To advance research in graph foundation models and facilitate scientific discovery in chemistry, materials science, and related fields, we will release the source code and pre-trained checkpoints.
1 Introduction
The deep learning revolution sparked by AlexNet (Krizhevsky et al., 2012) has driven remarkable progress in computer vision (CV) and natural language processing (NLP). The graph learning community similarly shifted from traditional machine learning to deep learning with the rise of graph neural networks (GNNs) (Kipf & Welling, 2017; Hamilton et al., 2017; Zhang & Chen, 2018; Wu et al., 2020).
Today, transformers dominate CV (Dosovitskiy et al., 2021; Liu et al., 2021) and NLP (Devlin et al., 2019; Radford et al., 2018), scaling to billions of parameters (Liu et al., 2021; Brown et al., 2020) and achieving superhuman performance on benchmarks like ImageNet (Deng et al., 2009) and GLUE (Wang et al., 2019). These advances underpin transformative applications such as ChatGPT (Open-AI, 2023) and Midjourney (Midjourney, 2023).
Despite progress, GNNs remain constrained by over-smoothing (Rusch et al., 2023) and over-squashing (Alon & Yahav, 2021), limiting their scalability and capacity to leverage large-scale graph data. Recent efforts to adapt transformers to graphs (Ying et al., 2021; Kim et al., 2022; Luo et al., 2023; Müller et al., 2023) show promise but face critical challenges: 1). Structural Bias: Most graph transformers (GTs) rely on handcrafted features or GNN modules to encode graph topology, compromising generalization. 2). Task Limitations: GTs excel at graph-level tasks but struggle with edge- and node-level objectives (Müller et al., 2023). 3). Pre-Training Gap: Unlike NLP’s success with self-supervised pre-training (Radford et al., 2018; Devlin et al., 2019), GTs lack effective frameworks for generative pre-training (Min et al., 2022; Müller et al., 2023).
In this work, we propose GraphGPT, a novel model for graph learning comprising three key innovations. 1). GET Backbone: A transformer-based architecture that operates on graph-equivalent token sequences via Eulerian paths, 2). Self-Supervised Pre-Training: Utilizing NTP and SMTP tasks (Radford et al., 2018; Chang et al., 2022), and 3). Task-Agnostic Fine-Tuning: Adapting the pre-trained model to supervised graph-, edge-, and node-level tasks.
Our contributions are summarized as follows:
Graph Eulerian Transformer (GET): We introduce GET, a novel architecture that leverages (semi-)Eulerian paths111For a quick recap, a graph with every node of even degree is Eulerian, and with exactly two nodes of odd degree is semi-Eulerian. to losslessly and reversibly convert graphs into token sequences. By integrating subgraph sampling and node identity encoding, GET efficiently processes graphs of arbitrary sizes. A standard transformer encoder or decoder is then applied to these sequences, eliminating the need for specialized architectural modifications.
Generative Pre-Training Framework: GraphGPT is pre-trained using NTP or SMTP tasks, offering three advantages: ) Captures structural and semantic graph patterns without handcrafted features or domain-specific architectures, ) Scales to over 2 billion parameters with sustained performance gains, and ) Enables effective graph generation through its sequence-based formulation.
Unified Task Formatting: We design a novel method to reformat graph-, edge-, and node-level tasks into sequences compatible with transformers. This approach allows downstream tasks to fully exploit pre-trained representations while unifying pretext and target task frameworks.
State-of-the-Art (SOTA) Performance: Extensive experiments on OGB datasets demonstrate GraphGPT’s superiority: it achieves SOTA results in graph- and edge-level tasks (e.g., molecular property prediction on PCQM4Mv2 and protein-protein interaction on ogbl-ppa), while delivering competitive performance in node-level tasks.
2 Approach
2.1 Overview
GraphGPT employs a three-stage framework: 1). Graph-to-Sequence Transformation of GET: The input graph is converted into a sequence of tokens via (semi-)Eulerian paths, ensuring a lossless, reversible mapping between the graph and its sequential representation. This transformation preserves node, edge, and attribute information while enabling compatibility with transformer architectures. 2). Self-Supervised Pre-Training: A standard transformer backbone (e.g., Llama; Touvron et al. (2023)) processes these sequences using the tasks NTP or SMTP (Radford et al., 2018; Chang et al., 2022). These tasks enable the model to learn structural and semantic graph patterns without task-specific supervision. 3). Task-Specific Fine-Tuning: The pre-trained model is adapted to downstream tasks—including graph classification/regression, link prediction, and node classification—by reformatting task objectives into sequence-based inputs. This unified approach maximizes the transfer of pre-trained knowledge.
2.2 Graph to Sequence of Tokens
To convert graphs into token sequences, we employ distinct strategies based on graph size:
Small graphs (e.g., molecular graphs) are directly serialized using the method in §2.2.1.
Large graphs (with up to billions of nodes/edges and rich attributes) are first decomposed into subgraphs via the sampling process described in §2.2.2. Node identity preservation (§2.2.3) ensures structural consistency during this decomposition. These subgraphs are then serialized using §2.2.1’s guidelines.
2.2.1 Serializing graphs with (semi-)Eulerian Paths
We propose a lossless, reversible graph serialization method based on traversing all edges and nodes via (semi-)Eulerian paths. This approach guarantees:
1.
Complete representation of nodes and edges in the sequence.
2.
Bijective mapping between the graph and its serialized form.
Algorithmic Foundation: The problem aligns with the Chinese Postman Problem (Mei-Ko, 1962; Edmonds & Johnson, 1973), which seeks the shortest path traversing all edges. For graphs lacking (semi-)Eulerian properties, we apply Eulerization (Edmonds & Johnson, 1973; Daubechies & Hughes, 2009), duplicating minimal edges to create an Eulerian multigraph.
Implementation Steps:
1.
Path Identification:
Check Eulerian properties using established criteria (West et al., 2001).
If non-Eulerian, perform Eulerization to enable path traversal.
Randomly sample one valid path from possible candidates, introducing stochasticity as a data augmentation strategy akin to computer vision techniques (Perez & Wang, 2017).
2.
Node Re-indexing:
Assign indices based on nodes’ first appearance in the path (e.g., Fig. 1: node , ).
Introduce cyclic re-indexing: , where is a random integer and (hyperparameter) exceeds the maximum node count. This ensures uniform token training by distributing index frequencies.
3.
Attribute Handling:
Discrete attributes: Direct tokenization.
Continuous attributes: Digit-wise tokenization (e.g., ).
Edge directionality: Distinct tokens for incoming/outgoing edges (e.g., , ).
Three attribute integration strategies (Fig. 1): short, long, and prolonged formats.
4.
Disconnected Components:
Connect components by adding synthetic edges between randomly selected nodes.
Label these edges with a dedicated EDGE_JUMP token and default attribute tokens.
Theoretical Guarantee: The serialization is lossless and reversible (up to isomorphism) per the Eulerian Path Theorem: reconstructing edges from adjacent tokens recovers the original graph structure (Grohe & Schweitzer, 2020). For example, in Fig. 1’s sequence, connecting consecutive tokens () reconstructs all edges, yielding a graph isomorphic to the input.
2.2.2 Subgraph Sampling
Directly serializing large graphs into sequences via the method in §2.2.1 produces excessively long sequences that exceed transformer context limits. While truncating such sequences is possible, this approach faces two critical issues:
1.
Computational Overhead: Eulerization and path identification for massive graphs are computationally prohibitive.
2.
Inconsistent Training: Sequence fragmentation introduces mismatches between pre-training and fine-tuning data formats.
To address these challenges, we adopt subgraph sampling—a scalable strategy that decomposes large graphs into smaller, manageable subgraphs before serialization.
Implementation:
We use the ShaDowKHop sampler (Zeng et al., 2021) to extract localized subgraphs centered on randomly selected nodes or edges.
Sampler parameters (e.g., hop depth, neighbor count) are preconfigured to ensure generated sequences fit within the transformer’s context window. These parameters are dataset- and hardware-dependent (see App. A.1 for configuration details).
2.2.3 Node Identity Encoding
Preserving global node identities during subgraph sampling is essential to avoid information loss. While unique token-based encoding (via learnable embeddings) is theoretically viable, it becomes impractical for graphs with billions of nodes due to:
Vocabulary Explosion: A 10-billion-node graph would require a vocabulary of size .
Memory Constraints: Corresponding embedding matrices become prohibitively large.
Solution: Multi-Token Node Encoding. We propose encoding each node as a combination of tokens, reducing vocabulary size exponentially. For example: A -node graph can be uniquely represented with two tokens from a -size vocabulary (). Graph partitioning via METIS (Karypis & Kumar, 1997) enables this by dividing the graph into clusters, each containing nodes.
Trade-offs: Increasing (e.g., ) allows smaller vocabularies () but lengthens sequences. This mirrors variable-length encodings like utf-8 (Allen et al., 2012, Chapter 2), where characters are represented by 1–4 bytes (vocabulary size ).
Our ablation studies (§3.5.2) demonstrate this method’s effectiveness in preserving node identity while maintaining tractable model scaling.
2.3 Modeling with the transformer decoder/encoder
We demonstrate how the transformer architecture processes graph token sequences under a unified pre-training and fine-tuning paradigm for diverse graph tasks.
2.3.1 Pre-training with the NTP or SMTP tasks
Self-supervised pre-training has proven critical for success in NLP (Devlin et al., 2019; Radford et al., 2019) and CV (He et al., 2022; Chang et al., 2022). We adopt two foundational generative tasks: NTP, which enables SOTA performance in NLP (Brown et al., 2020), and SMTP, which extends masked prediction with scheduled masking rates.
Implementation Details:
Masking (SMTP): For node-level masking, all occurrences of a masked node in the Eulerian sequence are hidden to prevent leakage (e.g., two occurrences of node in Fig. 1 are masked concurrently).
Mask Scheduling (SMTP): Following Chang et al. (2022), we use a linear scheduling function, which empirically balances training stability and performance.
2.3.2 Fine-tuning on downstream graph tasks
We adapt the pre-trained transformer to supervised tasks by reformatting the sequence-based inputs, ensuring alignment with pre-training:
1.
Graph-Level Tasks (e.g., classification/regression): Append a special GSUM token to the sequence.
2.
Edge-Level Tasks (e.g., link prediction): Append tokens of the target edge’s source and destination nodes.
3.
Node-Level Tasks (e.g., node classification): Append the target node’s token to the sequence.
The final token’s output is fed to a randomly initialized multilayer perceptron (MLP).
Parameter Update:
Transformer weights are initialized from pre-trained checkpoints.
MLP layers are trained from scratch.
All parameters are updated during fine-tuning.
3 Experiments
3.1 Datasets
Recent advances in AI for scientific discovery (Wang et al., 2023) motivate our evaluation of GraphGPT on large-scale scientific graph datasets spanning physics, chemistry, and bioinformatics. To demonstrate its versatility across graph tasks, we select benchmarks for graph-, edge-, and node-level objectives:
Graph-level: PCQM4Mv2 (quantum chemistry), ogbg-molpcba (molecular property prediction) and Triangles (triangles counting).
Edge-level: ogbl-ppa (protein-protein associations) and ogbl-citation2 (citation networks).
Node-level: ogbn-proteins (protein interaction networks) and ogbn-arxiv (paper categorization).
PCQM4Mv2 contains million organic molecules from PubChemQC (Nakata & Shimazaki, 2017). Nodes represent atoms (9D attributes: atomic number, chirality, etc.), and edges denote chemical bonds (3D attributes: bond type, stereochemistry, conjugation).
ogbg-molpcba is a smaller molecular dataset (Wu et al., 2017) with the same node/edge attributes.
Triangles (Knyazev et al., 2019) contains 45k graphs (no node/edge attributes).
ogbl-ppa: Nodes are proteins from 58 species; edges represent functional associations (Szklarczyk et al., 2019).
ogbl-citation2: A directed citation network with million papers (nodes) and million edges.
ogbn-proteins: Undirected, weighted graph of 132,534 proteins (nodes) with 8D edge attributes encoding association strengths.
ogbn-arxiv: Citation network of 169,343 papers; tasks involve predicting 40 subject categories.
Empirical results demonstrate that SMTP pre-training achieves superior or comparable performance across all benchmarks. Unless otherwise specified, reported GraphGPT results utilize SMTP pre-training.
3.2 Graph-level tasks
| Models | MAE | Params | |
| Valid | Test | ||
| GCN | 0.1379 | 0.1398 | 2.0M |
| GIN | 0.1195 | 0.1218 | 3.8M |
| GCN-VN | 0.1153 | 0.1152 | 4.9M |
| GIN-VN | 0.1083 | 0.1084 | 6.7M |
| TokenGT | 0.0910 | 0.0919 | 48.5M |
| Graphformer | 0.0864 | N/A | 48.3M |
| GPS-Deep | 0.0852 | 0.0862 | 138.1M |
| GPS++ (no 3D) | 0.0818 | N/A | 40.0M |
| GPTrans-L | 0.0809 | 0.0821 | 86.0M |
| GraphGPT-B12 | 0.0807 | N/A | 113.6M |
| GraphGPT-B24 | 0.0793 | N/A | 227.3M |
| GraphGPT-B48 | 0.0792 | 0.0804 | 453.4M |
We evaluate GraphGPT on two molecular datasets where tasks involve predicting quantum chemical properties solely from 2D molecular graphs—a practical alternative to relying on 3D equilibrium structures. Specifically: PCQM4Mv2 predicts the HOMO-LUMO energy gap, and ogbg-molpcba predicts 128 binary molecular properties.
On PCQM4Mv2, GraphGPT achieves a test MAE of 0.0804, significantly outperforming the previous SOTA (0.0821, Chen et al. (2023b)).
Compared to GTs like TokenGT, Graphformer, GPS, and GPTrans—which require handcrafted features or intricate architectures to encode structural information—GraphGPT attains superior performance without manual feature engineering. It also surpasses GNNs by a substantial margin (Table 1).
Analysis:
Lossless Serialization: The Eulerian path-based serialization and generative pre-training enable GraphGPT to fully capture structural and semantic graph information.
Scalability: While GTs with fewer parameters often plateau when scaled (Shi et al., 2022), GraphGPT shows consistent improvement up to 400M parameters.
Parameter Efficiency: GraphGPT’s larger parameter count may reflect its capacity to implicitly learn features that other GTs encode manually. Generative pre-training also allocates model capacity to generation tasks, potentially limiting discriminative performance at smaller scales.
Limitations: Pre-training with additional SMILES-formatted molecular data yielded diminishing returns, suggesting saturation in 2D structural information. Incorporating 3D molecular data could address this.
Transfer Learning: When fine-tuned on ogbg-molpcba, our PCQM4Mv2-pretrained model achieves results exceeding powerful GNNs (GCN, GIN) and matching SOTA GTs (Table 2).
| Models | Average Precision (%) | Params | |
| Test | Valid | ||
| GCN | 0.57M | ||
| GIN | 1.92M | ||
| GINE-VN | 6.1M | ||
| NGIN-VN | 44.19M | ||
| 3.84M | |||
| Graphormer-L† | 119.5M | ||
| EGT-Larger† | N/A | 110.8M | |
| GRPE-Large† | N/A | 118.3M | |
| GPTrans-L† | N/A | 86.0M | |
| GraphGPT-B | 113.6M | ||
| GraphGPT-B | 227.3M | ||
3.2.1 Graph Structure Understanding
To evaluate GraphGPT’s ability to learn structural patterns through generative pre-training, we use the Triangles dataset with the task of counting triangles. The dataset is split into: 1). Training/Validation: 30k and 5k small graphs ( nodes); 2). Testing: 5k small graphs (Test-small) and 5k large graphs (25–100 nodes, Test-large).
This task is challenging even for in-distribution (ID) graphs and considerably harder for out-of-distribution (OOD) graphs.
Pre-Training Setup: We augment pre-training with diverse datasets, i.e., Reddit-threads (Rozemberczki et al., 2020), Erdős-Rényi random graphs (Erdos et al., 1960), and Internal real-world graphs (See Table 8, Appendix A).
Key Observations (Table 3):
Pre-Training Efficacy: GraphGPT achieves comparable accuracy to GTs on ID graphs and superior OOD generalization (lower variance). This demonstrates that generative pre-training effectively encodes structural knowledge transferable to downstream tasks.
Impact of Graph Types: Pre-training on real-world graphs (e.g., internal datasets) outperforms random Erdős-Rényi graphs, suggesting meaningful structural patterns in real-world data enhance model learning.
Dataset Diversity: Combining Triangles with diverse datasets (Reddit-threads, internal graphs) yields better performance than pre-training on Triangles alone. This highlights the importance of diverse pre-training data for learning generalizable structural patterns.
Attributed Graphs: Models pre-trained on attributed graphs (PCQM4Mv2, ogbl-ppa, ogbn-proteins) and fine-tuned on Triangles achieve significant improvements: vs. (baseline GET without pre-training). This confirms that structural knowledge is obtained even when pre-training includes node/edge attributes.
| Models | Accuracy (%) | Params | |
| T-small | T-large | ||
| GIN | 0.15M | ||
| Transformer | 0.2M | ||
| Transformer-LapPE | 0.2M | ||
| Transformer-RWSE | 0.2M | ||
| Graphormer | 0.2M | ||
| GET-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-B | 113.5M | ||
| GraphGPT-M | 33.7M | ||
| GraphGPT-B | 113.5M | ||
3.3 Edge-level tasks
We evaluate GraphGPT on link prediction using the ogbl-ppa and ogbl-citation2 datasets. Results are summarized in Table 4.
Performance Superiority: GraphGPT significantly outperforms all baseline methods, including GNNs, heuristic models, and latent-factor approaches, across both datasets. This underscores the effectiveness of generative pre-training and sequence-based modeling for edge-level tasks.
Scalability: GraphGPT scales seamlessly to 2 billion parameters, achieving sustained performance gains with increasing model size. This motivates future exploration of even larger architectures and datasets.
Transformer Efficacy: To our knowledge, GraphGPT is the first transformer-based model to achieve SOTA results on ogbl-ppa and ogbl-citation2, demonstrating the viability of sequence-driven architectures for large-scale edge-level tasks.
| Models | ogbl-ppa | ogbl-citation2 |
| HR@100 (%) | MRR (%) | |
| Common Neighbor | ||
| Adamic Adar | ||
| Resource Allocation | ||
| Node2Vec | ||
| Matrix Factorization | ||
| GCN | ||
| GraphSAGE | ||
| SEAL | ||
| AGDN | ||
| SIEG | ||
| MPLP | ||
| RefinedGAE | ||
| GraphGPT-M | ||
| GraphGPT-B | ||
| GraphGPT-XXL | N/A |
3.4 Node-level tasks
We evaluate GraphGPT on two node-level benchmarks: ogbn-proteins predicts 112 binary protein function labels, and ogbn-arxiv classifies arXiv papers into 40 subject categories. Results are summarized in Table 5.
ogbn-proteins: GraphGPT surpasses well-tuned GNN baselines (GCN, GraphSAGE, GAT) and significantly outperforms graph transformers (GTs). Remarkably, GraphGPT achieves competitive performance with input subgraphs of nodes, while SOTA GNNs like AGDN (Sun et al., 2020) require subgraphs with nodes.
ogbn-arxiv: GraphGPT delivers performance comparable to or approaching SOTA graph transformers and optimized GNNs.
The strong performance with minimal neighborhood sampling suggests that generative pre-training effectively encodes global structural and semantic graph information into node token embeddings and transformer parameters. This contrasts with traditional GNNs, which rely on extensive local aggregation for feature propagation.
| Models | ogbn-proteins | ogbn-arxiv |
| ROC-AUC (%) | Accuracy (%) | |
| GCN | ||
| GraphSAGE | ||
| GAT | ||
| DRGAT | N/A | |
| AGDN | ||
| DeeperGCN | ||
| GraphGPS | ||
| NAGphormer | ||
| Exphormer | ||
| GOAT | ||
| NodeFormer | ||
| SGFormer | ||
| Polynormer | ||
| GraphGPT-S | ||
| GraphGPT-M | ||
| GraphGPT-B |
3.5 Ablation Study
We analyze the impact of two core components of GraphGPT: pre-training and node identity encoding.
3.5.1 Pre-training
The self-supervised NTP or SMTP tasks are central to GraphGPT’s success. As shown in Table 6, pre-training delivers performance improvements of 10–100% across graph-, edge-, and node-level tasks. These gains highlight its role in enabling the model to learn intrinsic graph structural patterns and capture semantic relationships inherent in node and edge attributes.
| Datasets | Pre-training | Test | Valid |
| PCQM4Mv2 | ✗D | N/A | 0.0978 |
| ✗E | N/A | 0.0856 | |
| NTP | N/A | 0.0875 | |
| SMTP | N/A | 0.0807 | |
| ogbg-molpcba | ✗D | 12.80 | 13.31 |
| ✗E | 25.80 | 26.33 | |
| NTP | 23.85 | 27.77 | |
| SMTP | 27.56 | 28.74 | |
| SMTP∗ | 27.20 | 28.49 | |
| SMTP∗ + FT† | 28.07 | 29.01 | |
| ogbl-ppa | ✗D | 41.28 | 40.14 |
| ✗E | 42.13 | 41.57 | |
| NTP | 55.56 | 54.87 | |
| SMTP | 55.68 | 54.93 | |
| ogbn-proteins | ✗D | 57.52 | 61.19 |
| ✗E | 53.20 | 56.39 | |
| NTP | 75.61 | 80.47 | |
| SMTP | 83.56 | 87.73 |
3.5.2 Node Identity Encoding
Node identity encoding (see §2.2.3)—representing nodes in large graphs as unique multi-token sequences—is critical for edge- and node-level tasks. Using GraphGPT-mini (a lightweight variant to conserve computational resources), we demonstrate that this method significantly enhances performance (Table 7). Further implementation details are provided in Appendices F and A.
| Datasets | Params | NIE | Test | Valid |
| ogbl-ppa | 14.75M | ✗ | 44.38 | 45.08 |
| ✓ | 55.56 | 54.87 | ||
| ogbn-proteins | 10.76M | ✗ | 60.22 | 65.66 |
| ✓ | 75.61 | 80.47 |
4 Limitations
We critically assess the limitations of GraphGPT to contextualize its applicability and inspire future improvements.
Transferability. GraphGPT’s reliance on dataset-specific pre-training limits its ability to generalize across domains with divergent semantics (e.g., social networks vs. molecular graphs). However, it demonstrates robust cross-dataset structural understanding (§3.2.1) and effective intra-domain transferability, as evidenced by molecular data experiments (§3.2).
Dataset size. Performance on some small- to medium-sized datasets (e.g., ogbn-arxiv) lags behind traditional GNNs. This can be mitigated by expanding datasets with semantically aligned data.
Computational Cost. Pre-training on large-scale graphs (ogbn-proteins, ogbl-ppa) or extensive small graphs (PCQM4Mv2) with 50M+ parameters is resource-intensive. For example, pre-training GraphGPT-B (100M+ parameters) on PCQM4Mv2 with tokens requires V100 GPU hours, and fine-tuning incurs V100 GPU hours per epoch.
While GraphGPT is less practical for small datasets due to compute-performance trade-offs, it excels with large-scale data. Emerging techniques like quantization (Dettmers et al., 2022; Frantar et al., 2022), distributed training frameworks (Rasley et al., 2020; Shoeybi et al., 2019), and transformer optimizations (Dao, 2023) are poised to alleviate these costs.
Future Directions. These limitations highlight opportunities for research in cross-domain transfer, data-efficient training, and scalable architectures.
5 Related works
Graph Neural Networks (GNNs) GNNs have dominated graph learning for decades, with numerous variants achieving strong performance across tasks (Wu et al., 2020). However, they face fundamental limitations such as over-smoothing and over-squashing (Rusch et al., 2023; Alon & Yahav, 2021), which hinder their scalability and ability to model long-range dependencies.
Graph Transformers (GTs) Inspired by transformers’ success in NLP and CV, recent work has adapted these architectures to graphs (Ying et al., 2021; Rampásek et al., 2022; Müller et al., 2023). While GTs achieve competitive results on large-scale graph-level tasks (Müller et al., 2023), they typically rely on handcrafted structural features or GNN-based modules to encode graph topology—either in input representations (Ying et al., 2021; Kim et al., 2022; Masters et al., 2023) or attention mechanisms (Ying et al., 2021; Chen et al., 2022; Luo et al., 2023).
Pre-training and fine-tuning The self-supervised pre-training and supervised fine-tuning paradigm, popularized by transformers (Vaswani et al., 2017), revolutionized NLP (Devlin et al., 2019; Radford et al., 2018). Scaling this approach with web-scale data (Brown et al., 2020) and techniques like instruction tuning (Wei et al., 2021) or reinforcement learning from human feedback (Ouyang et al., 2022) further advanced the field. In CV, self-supervised methods like MAE He et al. (2022) and MaskGIT (Chang et al., 2022)demonstrated that masked prediction tasks (e.g., reconstructing masked image patches) enable transformers to achieve SOTA results.
6 Conclusions
We introduce GraphGPT, a novel model built on the GET backbone, which achieves SOTA or near-SOTA performance across graph-, edge-, and node-level tasks on large-scale benchmarks. By unifying pretext and downstream tasks into a sequence-based paradigm, GraphGPT demonstrates strong transferability in capturing both structural graph patterns and domain-specific knowledge (e.g., molecular properties). Notably, scaling GraphGPT to billions of parameters yields consistent performance gains, highlighting its potential as a foundation for graph-centric foundation models.
Looking ahead, GraphGPT’s architecture is inherently scalable—capable of expanding to hundreds of billions of parameters—and offers promising avenues for integration or alignment with large language models (LLMs), bridging graph reasoning and textual intelligence.
Appendix A Datasets
The detailed statistics of the datasets are in the Tab. 8.
| datasets | # of graphs | avg # of nodes | avg # of edges | task-type | metrics |
| PCQM4Mv2 | 3,746,619 | 14.14 | 14.56 | regression | MAE |
| ogbg-molpcba | 437,929 | 26.0 | 28.1 | multi-label BC | AP |
| reddit-threads | 203,088 | 23.9 | 24.9 | BC | ROC-AUC |
| Triangles | 45,000 | 20.9 | 32.7 | multi-class classification | ACC |
| Internal dataset | 3,100,000 | 24.8 | 54.7 | N/A | N/A |
| Random Graphp=0.03 | 3,100,000 | 67.1 | 74.8 | N/A | N/A |
| ogbl-ppa | 1 | 576,289 | 30,326,273 | BC | HR@100 |
| ogbl-citation2 | 1 | 2,927,963 | 30,561,187 | BC | MRR |
| ogbn-proteins | 1 | 132,534 | 39,561,252 | multi-label BC | ROC-AUC |
| ogbn-arxiv | 1 | 169,343 | 1,166,243 | multi-class classification | ACC |
A.1 Subgraph sampling
The subgraph sampling configurations for different datasets of large graphs are shown in the Tab. 9.
| dataset | sampling | depth () | # neighbors () | seq-len |
| ogbl-ppa | edge-ego | 1 | 14 | 90 |
| 1 | 30 | 280 | ||
| ogbl-citation2 | edge-ego | 1 | 14 | 60 |
| 1 | 20 | 90 | ||
| 1 | 30 | 130 | ||
| ogbn-proteins | node-ego | 9 | 1 | 20 |
| 20 | 1 | 50 | ||
| 40 | 1 | 120 | ||
| 60 | 1 | 200 | ||
| ogbn-arxiv | node-ego | 1 | 30 | 30 |
| 1 | 40 | 40 |
Appendix B Models
We list the model specifics in the Tab. 10. We experiment with eight different scales of models.
| Model-size | Hidden-size | # of layers | # of heads | Params (excluding embed) |
| Mini | 256 | 4 | 4 | 4.2M |
| S (Small) | 512 | 4 | 8 | 16.8M |
| M (Medium) | 512 | 8 | 8 | 33.6M |
| B / B12 (Base) | 768 | 12 | 12 | 113.2M |
| B24 (Base24) | 768 | 24 | 12 | 226.5M |
| B48 (Base48) | 768 | 48 | 12 | 453.0M |
| L (Large) | 1024 | 24 | 16 | 402.7M |
| XXL (XXLarge) | 1600 | 48 | 25 | 2.0B |
Appendix C Implementation details
C.1 Graphs to Sequences of tokens
The code is implemented with Pytorch. We use torch-geometric (Fey & Lenssen, 2019) to preprocess graphs, e.g., subgraph sampling and etc. We use Networkx (Hagberg et al., 2008) to Eulerize the (sub)graphs if needed and then find the (semi-)Eulerian paths. We write our own tokenizer to convert the (semi-)Eulerian paths to sequences of tokens. The token vocabulary is built for each dataset separately.
C.2 Model backbone
We adopt Llama’s transformer architecture (Touvron et al., 2023) implemented in the Hugging Face’s transformers package (Wolf et al., 2020) as our backbone for the NTP pre-training. To turn it into an encoder for the SMTP pre-training, we change the causal attention mask to bi-directional attention mask. We do not use Llama’s pre-trained weights. Instead, we train models of different scales as in Tab. 10 with random parameters initialization.
C.3 Training
The models are pre-trained and fine-tuned on GPU clusters of A800-80G using DeepSpeed’s stage-2 schema with mixed precision (Rasley et al., 2020) or bf16. We use AdamW (Loshchilov & Hutter, 2017) optimizer with learning rate scheduler. To make full use of computing power, we pack several graph sequences together into one entry to maximize the context window (Raffel et al., 2020). The detailed configurations for each dataset are set separately and can be found in the corresponding sections in the following.
C.4 Vocabulary
In NLP, the vocabulary is usually built by tokenizing the text data with the byte-pair encoding (BPE) algorithm (Sennrich et al., 2016). The resulting unique tokens form the vocabulary, and they are usually frequent subwords in the text corpus.
In our GraphGPT, the vocabulary is constructed very differently. We split the vocabulary into two parts. The first part contains the structural and special tokens, and they are dataset agnostic and transferable among datasets. The second part consists of tokens that encode the semantics information of the dataset, such as node and edge attributes.
Appendix D Graphs to sequences of tokens
In this section, we show some examples of turning graphs to sequences of tokens.
D.1 Molecular graphs to tokens
Below is one example of 2D molecular graphs in the ogbg-molpcba dataset in torch-geometric data format (Fey & Lenssen, 2019).
The graph has 4 nodes and 3 edges. The source and destination nodes of the edges are recorded in ‘edge_index’, and its dimension is () for undirected graphs. ‘x’ is the node attributes of 9 dimensions, and ‘edge_attr’ stores the edge attributes of 3 dimensions.
The node and edge attributes of the graphs are numbers. If we directly discretize them into tokens, i.e., using one token to represent each unique number, the numbers that appear few times in the dataset cannot be well-trained. At the same time, the vocabulary may blow up. Therefore, we split them into single digits and represent them with the combination of the following tokens. They are dataset agnostic, and can be shared across different datasets.
The resulting vocabulary is 556 for both ogbg-molpcba and PCQM4Mv2.
Below shows the tokens from one of the possible (semi-)Eulerain paths of the above molecular graph.
In the sequence of tokens above, for the node ‘1’, we can deduce that its 9 dimensional attributes are . Node ‘1’ is connected to ‘2’ with edge attributes . We set 0 as the the default value of the attributes in this dataset, and do not encode it into tokens.
In the (semi-)Eulerian path, a node may appear several times. We append its attributes tokens to one of its appearances randomly. This can prevent the model from copying the attributes from the previous appearance, and also shorten the resulting sequence.
For a graph obtained from Eulerization, an edge may present several times in the path. We apply the same logic to insert the edge attributes tokens.
As in the above sequence, node ‘2’ appears two times, and its node attributes tokens are appended after its second appearance. There is no tokens encode the edge attributes of edge between ‘2’ and ‘3’, which implies the edge attributes are default value .
D.2 Subgraphs to tokens
In edge/node-level tasks, we usually have one big graph. In this section, we use ogbl-ppa and ogbn-proteins datasets to show how to sample subgraphs from the big graph, and then transform the subgraph to sequences of tokens.
The whole ogbl-ppa dataset is summarized in torch-geometric format as follows.
It has 576289 nodes and 21231931 edges in the training data. ‘x’ is the one-hot representation of the species that the node (protein) belongs to.
We sample a subgraph from it as below.
It has 30 nodes, 42 edges as in ‘edge_index’. ‘x’ is the node attributes of 2 dimensions, and it encodes the node identity as described in Sec. 2.2.3. We partition the nodes (proteins) based on the associated species. The number of proteins inside each species varies from 616 to 41017. Finally we use 58 tokens for species and 41017 tokens for the local indices. Combined with the tokens for the structure and the special tokens, the total vocabulary is 41231.
Here ‘root_n_id’ records the two seed nodes, and the subgraph is sampled centered around them. The resulting tokens from one of the possible (semi-)Eulerian paths are:
In the ablation study on node identity encoding in Sec. 3.5.2, an example of the subgraph sampled from ogbl-ppa without identity encoding is shown below.
Different from the subgraph with node identity encoded in ‘x’, its node attribute ‘x’ contains only the information of the node’s (protein) hosting species. It cannot be used to uniquely identify the nodes. The vocabulary decreases from 41231 to 214.
The resulting tokens from one of its possible (semi-)Eulerian paths is below.
In the following, we use the ogbn-proteins dataset as the example. The entire dataset is a large graph as below.
It has 132,534 nodes and 39,561,252 edges. ‘node_species’ stores the species’ numeric id that the node (proteins) belongs to.
One sampled subgraph in the torch-geometric data format is:
It has 10 nodes, 11 edges as in ‘edge_index’. Edge attributes is stored in ‘edge_attr’ of dimension 8. ‘x’ is the node attributes of 2 dimensions, and it encodes the node identity as described in Sec. 2.2.3. Its first dimension (token) represents the species, and the second is local numbering of each protein inside its species. Similar to the ogbl-ppa dataset, the identity encoding of 132,534 nodes occupies 25,465 tokens in the vocabulary, and the total vocabulary is 25,620.
‘y’ records the labels for the supervised node-level task. ‘root_n_id’ represents the target node, and the subgraph is sampled centered around it.
The resulting tokens from one of the possible (semi-)Eulerian paths are as follows.
The original edge attributes are 8-dimensional vector of 3 decimal numbers from to . We split them into single digits and represent them with the combination of the digits tokens as in App. E.
To reduce the number of tokens in the resultant sequences further, we multiply the number with 1000 and then minus it by 1. So we do not need to encode ‘.’ any more. At the same time, we treat the value ( after the above transformation) as the default value and do not encode it with tokens.
Appendix E Graph-level task
E.1 PCQM4M-v2
The pre-training and fine-tuning configurations for PCQM4M-v2 are in Tab. 11.
| pre-train | fine-tune | |
| model-size | Mini Small Medium Base Base24 Base48 | |
| batch-size | 1024 | |
| total | tokens | 32 epochs |
| warmup | tokens | 9.6 epochs |
| lr scheduler | Warmup & linear decay | Warmup & cosine decay |
| max-lr | ||
| min-lr | 0 | automatic set |
| Adam-betas | ||
| Adam-eps | ||
| max-grad-norm | 5 | 1 |
| weight-decay | 0.1 | 0.02 |
| attention-dropout | 0.1 | |
| path-dropout | 0 | |
| embed-dropout | 0 | |
| mlp-dropout | 0 | |
| LSI-val | NA | 1 |
| EMA | NA | |
| hidden-act | gelu | |
| MPE | 1024 | |
| TWE | FALSE | NA |
Appendix F Edge-level task
F.1 ogbl-ppa
We use two tokens for the node identity encoding introduced in Sec. 2.2.3. Specifically, we use the species to partition the nodes, so the first token represents the species, and the second is the local indices of proteins inside each species.
| pre-train | fine-tune | |
| model-size | Mini Small Medium Base Large XXLarge | |
| batch-size | ||
| total | tokens | epochs |
| warmup | tokens | epochs |
| lr scheduler | Warmup & linear decay | Warmup & cosine decay |
| max-lr | ||
| min-lr | automatic set | |
| Adam-betas | ||
| Adam-eps | ||
| max-grad-norm | 5 | 1 |
| weight-decay | 0.1 | 0 |
| attention-dropout | 0.1 | |
| path-dropout | ||
| embed-dropout | 0 | |
| mlp-dropout | 0 | |
| LSI-val | NA | |
| EMA | NA | |
| hidden-act | gelu | |
| MPE | ||
| TWE | FALSE | NA |
The pre-training and fine-tuning configurations for ogbl-ppa are listed in Tab. 12. The loss of pre-training versus the number of tokens is shown in Fig. 3.
The fine-tuning data consists of subgraphs induced by the positive edges for training and equal negative edges randomly sampled.
In general, a larger model results in lower pre-training loss, and better results in down-stream fine-tuning tasks.
F.2 ogbl-citation2
| pre-train | fine-tune | |
| model-size | Medium Base | |
| batch-size | ||
| total | tokens | epochs |
| warmup | tokens | epochs |
| lr scheduler | Warmup & linear decay | Warmup & cosine decay |
| max-lr | ||
| min-lr | automatic set | |
| Adam-betas | ||
| Adam-eps | ||
| max-grad-norm | 1 | |
| weight-decay | 0.1 | 0 |
| attention-dropout | 0.1 | |
| path-dropout | ||
| embed-dropout | 0 | |
| mlp-dropout | 0 | |
| LSI-val | N/A | |
| EMA | N/A | |
| hidden-act | gelu | |
| MPE | ||
| TWE | FALSE | N/A |
The pre-training and fine-tuning configurations for ogbl-citation2 are listed in Tab. 13. The subgraph sampling is edge-ego with and as in Tab. 9. The pre-training losses of the ogbl-citation2 dataset with different model sizes and subgraph sampling settings are shown in Fig. 4.
Appendix G Node-level task
G.1 ogbn-proteins
| pre-train | fine-tune | |
| model-size | Small Medium Base | |
| batch-size | ||
| total | tokens | epochs |
| warmup | tokens | epochs |
| lr scheduler | Warmup & linear decay | Warmup & cosine decay |
| max-lr | ||
| min-lr | automatic set | |
| Adam-betas | ||
| Adam-eps | ||
| max-grad-norm | 1 | |
| weight-decay | 0.1 | 0 |
| attention-dropout | 0.1 | |
| path-dropout | 0 | |
| embed-dropout | 0 | |
| mlp-dropout | 0 | |
| LSI-val | N/A | |
| EMA | N/A | |
| hidden-act | gelu | |
| MPE | ||
| TWE | FALSE | N/A |
G.2 ogbn-arxiv
| pre-train | fine-tune | |
| model-size | Small Medium Base | |
| batch-size | ||
| total | tokens | epochs |
| warmup | tokens | epochs |
| lr scheduler | Warmup & linear decay | Warmup & cosine decay |
| max-lr | ||
| min-lr | automatic set | |
| Adam-betas | ||
| Adam-eps | ||
| max-grad-norm | 1 | |
| weight-decay | 0.1 | 0 |
| attention-dropout | 0.1 | |
| path-dropout | 0 | |
| embed-dropout | 0 | |
| mlp-dropout | 0 | |
| LSI-val | N/A | |
| EMA | N/A | |
| hidden-act | gelu | |
| MPE | ||
| TWE | FALSE | N/A |